The probability that the sample man should not differ by population mean by more than a/2 units is 0.95, where σ is the population standard deviation. For the validity of the statement, what should be the size of the sample?

Select one:

a. 18

b. 16

c. 12

d. 20


1
Expert's answer
2021-10-28T11:48:44-0400

(Chebyshev's Theorem) The probability that any random variable "X" will assume a value within "k" standard deviations of the mean is at least "1 \u2014 1\/k^2."

That is, "P(\mu-k\sigma<X<\mu+k\sigma)\geq1-1/k^2."

Then


"1-1\/k^2=0.95=>1\/k^2=1\/20=>n=k^2=20"

d. 20